Projectile motion report:
Please help me answers (a) through (h)
Calculation Calculate vo from equation (8) g 980 cm/sec 2 1 joule 107 crgs 9 R CONCLUSION Summary Parts & q 2 Initial velocity of projectile Part II) vo cm/sec 20, 3 Initial velocity of projectile (Part I) Velocity v of system after impact (Part I), v 1312pm/sec 3HPg ec Kinetic energy of the system just after impact KE after KE after Kinetic Joules energy of the projectile just before impact, KE before ergs before joules Fractional change of Kinetic energy R Mass ratio 0.79 739 (a) which initial velocity, v o (Part I or Part is likely to be closer to the true value (b) Find the difference between your two determinations of v o (e) Why should your pendulum platform be clamped (d) Use data obtained in this lab to check whether momentum was conserved see equation (H), Part 1. e) Explain the loss of kinetic energy (if any) 00 Will friction in the pendulum (catch mechanlsm bearings cause your calculated velocity to be less than or greater than the velocity g) What effect does the force of gravity have on the horizontal velocity of the projectile (b) What effect does air resistance have on the range of the projectile
Answer #1

[a]

Initial velocity of projectile, Vo= 0.3743 m/sec is closer ot the true value [ true value= 0.3627m/s] .

[b]

Percentage difference of Initial velocity of projectile, Vo= 0.3743 m/sec is

% difference=[true value- Vo ]/ true value * 100%

% difference= [0.3627-0.3743]/ 0.3627 *100%= 3.2 %

Percentage difference of Initial velocity of projectile, Vo= 0.3743 m/sec is 3.2 %

Percentage difference of Initial velocity of projectile, Vo= 6.6277 m/sec is

% difference=[true value- Vo ]/ true value * 100%

% difference= [0.3627-6.6277 ]/ 0.3627 *100%=1727.3 %

Percentage difference of Initial velocity of projectile, Vo= 6.6277 m/sec is 1727.3 %

[c]

The pendulum will be held by a clamp stand to attain the experimental set up in equilbrium.

[d]

Conservation of momentum says that momentum before impact is equal to momentum after impact.

Pbefore= P after

mVo= [m+M] V

Here in both elastic and inelastic case, momentum is conserved,

[e]

Loss of Ke= KE after impact- KE before impact

Loss of Ke= = 1/2 [ m+M] V2- 1/2mVo2

[f]

Yes, Friction in the pendulum causes the calculated velocity to be less than the actual one.